# ViG-Bias: Visually Grounded Bias Discovery and Mitigation

### v1

@@ 4.3 Improving bias mitigation via visual explanations
where $n_{g}$ is the number of samples per group. The original B2T proposes inferring group annotations using the language-based zero-shot classifier, CLIP[19], and then re-train a classifier using the GroupDRO algorithm with the new group annotations. Here we evaluate the performance of GroupDRO when trained using the keywrods inferred by the original B2T and our proposed ViG-B2T.
@@ 6.2 Evaluating Bias Mitigation
We then evaluate the effectiveness of the identified slices and keywords in reducing spurious correlations. Initially, we assess the average and worst-group accuracies for zero-shot classification tasks using various prompting strategies. Then, we compare debiasing models using the GroupDRO algorithm, each using the discovered keywords to infer the new group annotations across all datasets.
@@ 6.3 Ablations
We perform ablation studies considering an alternative visual explanation method (ScoreCAM [26], which differently from GradCAM, does not rely on gradients to construct the class activation maps) and different thresholds $\tau$ for the generation of the binary masks. Figure 5 shows results both for bias discovery (Figure 5(a)) and mitigation (Figure 5(b)). We aim to assess the stability of the visual explanation methods across different threshold levels. We find that both, GradCAM and ScoreCAM achieve approximately similar worst group accuracies for different values of $\tau$. For instance, ScoreCAM gets good worst-group accuracy at lower threshold values, whereas GradCAM [21] requires increasing threshold values. Nevertheless, the worst group accuracy remains approximately constant for both methods. Finally, some qualititave results from these visualizations are provided in Figure 6 for different attention methods. Some more visual results are also included in Supplementary Material C.
@@
Figure 5: Ablating the GradCAM [21] mask threshold $\tau$ on the CelebA dataset for bias discovery (a) and mitigation (b).
Figure 6: Visual explanations of other CAM-variants (GradCAM++ [4], Fullgrad [23] and ScoreCAM [26]) for a biased model trained on CelebA. Compared to other CAM methods, ScoreCAM tends to better highlight the region related to the spurious feature (in this case, instead of focusing on the hair, it points to the face), for images from groups where the spurious correlation holds.
@@ 7 Conclusion
Bias discovery and mitigation remain critical challenges in the deployment of fair and reliable machine learning models. In this paper, we introduced Visually Grounded Bias Discovery (ViG-Bias), a novel method that leverages visual explanations to uncover and address biases in visual recognition models. By integrating multimodal embeddings with visual explanation techniques, ViG-Bias not only identifies biases with higher precision but also provides insights into the nature of these biases, facilitating the development of more equitable models.
Our approach represents a significant step forward in unsupervised bias discovery and mitigation research by offering a method that is both effective and interpretable, while easily adaptable to multiple existing frameworks. The systematic improvements observed across various datasets underscore the utility of visual explanations in enhancing model fairness and reliability, by reducing their dependence on spurious correlations.
As future work, we aim to investigate the integration of ViG-Bias with other bias detection and mitigation frameworks to further enhance its effectiveness and adaptability.
8 Acknowledgments
We gratefully acknowledge the DATAIA program for supporting EF as a visiting professor at Université Paris-Saclay, NVIDIA Corporation for provinding GPU computing, the support of Agencia Nacional de Promoción de la Investigación, el Desarrollo Tecnológico y la Innovación (Argentina), the Google Award for Inclusion Research (AIR) Program and the SticAmSud program.
Appendix 0.A Datasets
@@ Appendix 0.B Implementation Details
Appendix 0.C Additional qualitative visualizations and results
0.C.1 Discovered Keywords by ViG-B2T
@@
0.C.2 Complete Tables of Keywords
Tables 5(b)-8(b) present a comparison between the keywords extracted by B2T and ViG-B2T. We can observe that ViG-B2T extracted keywords which are more relevant to the problem as they are related to spurious correlations, effectively highlighting biases in the dataset. Specifically, ViG-B2T extracted keywords which are related to gender (male or female), whereas B2T’s keywords only mention person without specifying gender, which is crucial information for bias discovery in this case.
@@ 0.C.2 Complete Tables of Keywords
0.C.3 Comparing biased and de-biased models using GradCAM
@@
0.C.4 Slices discovered by ViG-FACTS

### v3

@@ 4.3 Improving bias mitigation via visual explanations
where $n_{g}$ is the number of samples per group. The original B2T proposes inferring group annotations using the language-based zero-shot classifier, CLIP[19], and then re-train a classifier using the GroupDRO algorithm with the new group annotations. Here, we evaluate the performance of GroupDRO when trained using the keywords inferred by the original B2T and our proposed ViG-B2T.
@@ 6.2 Evaluating Bias Mitigation
We then assess the effectiveness of identified slices and keywords in reducing spurious correlations. Initially, we assess the average and worst-group accuracies for zero-shot classification using various prompting strategies. Then, we compare debiasing models using the GroupDRO algorithm, each using the discovered keywords to infer the new group annotations across all datasets.
@@ 6.3 Ablations
We considered an alternative VE method (ScoreCAM [26], which differently from GradCAM, does not rely on gradients to construct the class activation maps) and different thresholds $\tau$ for the generation of the binary masks. Figure 5 shows results both for bias discovery (Figure 5(a)) and mitigation (Figure 5(b)). We aim to assess the stability of the visual explanation methods across different threshold levels. We find that both, GradCAM and ScoreCAM achieve approximately similar worst group accuracies for different values of $\tau$. For instance, ScoreCAM gets good worst-group accuracy at lower threshold values, whereas GradCAM [21] requires increasing threshold values. Nevertheless, the worst group accuracy remains approximately constant for both methods. Finally, some qualititave results from these visualizations are provided in Figure 6 for different attention methods. Additional results are also included in Supplementary Material C.
@@
Figure 5: Ablating the GradCAM [21] mask threshold $\tau$ on the CelebA dataset for bias discovery (a) and mitigation (b). The main hyperparameter that is important for our model is the parameter $\tau$ that integrates the optimal threshold for the mask. We optimize this parameter depending on the dataset, however, our method performs better than the baseline in almost all the different choices of $\tau$. Empirically, as a rule of thumb, setting $\tau=0.5$ is a good starting point.
Figure 6: Visual explanations of other CAM-variants (GradCAM [21], GradCAM++ [4], Fullgrad [23] and ScoreCAM [26]) for a biased model trained on CelebA. Compared to other CAM methods, ScoreCAM tends to better highlight the region related to the spurious feature (in this case, instead of focusing on the hair, it points to the face), for images from groups where the spurious correlation holds.
@@ 7 Conclusion
In this paper, we introduced Visually Grounded Bias Discovery (ViG-Bias), a novel method that leverages visual explanations to uncover and address biases in visual recognition models. By integrating multimodal embeddings with visual explanation techniques, ViG-Bias not only identifies biases with higher precision but also provides insights into the nature of these biases. Our approach represents a significant step forward offering a method that is both effective and interpretable, while easily adaptable to multiple existing frameworks. However, it is not free from limitations. The type of spurious correlation that can be discovered by ViG-Bias is limited to those that can be attributed by VEs. For example, spurious correlations not related to localized artifacts or objects, like global intensity distribution shifts due to noise or different acquisition devices, will not be detected by ViG-Bias. As for Societal Impact, we believe it is overall positive as our method improves bias detection and mitigation, ultimately contributing to more fair AI systems. As future work, we aim to investigate the integration of ViG-Bias with other bias detection and mitigation frameworks to further enhance its effectiveness and adaptability.
Acknowledgments
We gratefully acknowledge the DATAIA program for supporting EF as a visiting professor at Université Paris-Saclay, NVIDIA Corporation for provinding GPU computing, the support of Agencia Nacional de Promoción de la Investigación, el Desarrollo Tecnológico y la Innovación (Argentina), the Google Award for Inclusion Research (AIR) Program and the SticAmSud program. This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program as well as the ANR Hagnodice ANR-21-CE45-0007.
Appendix 0.A Appendix
0.A.1 Datasets
@@ 0.A.2 Implementation Details
0.A.3 Additional results
0.A.3.1 Dependecy of Vig-Bias on Visual Explanation Methods
In this section we include additional results for the NICO++ dataset, obtained using visual explanations (VEs) different from GradCAM and ScoreCAM, to demonstrate that our method can be used in combination with different VEs. Tab. 5 presents its effectiveness in other variants of CAM. As can be observed, all methods perform better than the baseline (and some even better than the initial GradCAM), highlighting the potential of this simple idea and opening the door for future research where more powerful explanation methods could be incorporated.
Table 5: Ablating VEs on the NICO++95. Percentages indicate the Precision@k metric (same as Tab. 1 in the main manuscript).
Method VE methods
GradCAM GradCAM++ FullGrad ScoreCAM
DOMINO $24.0\%$ $24.0\%$ $24.0\%$ $24.0\%$
ViG-DOMINO (ours) $\mathbf{25.4\%}$ $\mathbf{32.2\%}$ $\mathbf{30.7\%}$ $\mathbf{26.5\%}$
FACTS $60.8\%$ $60.8\%$ $60.8\%$ $60.8\%$
ViG-FACTS (ours) $\mathbf{65.0\%}$ $\mathbf{64.8\%}$ $\mathbf{66.6\%}$ $\mathbf{65.2\%}$
0.A.3.2 Correlation between visual explanations and spurious features.
Here we include an additional experiment to analyze how different visual explanation methods tend to highlight spurious features when computed on biased models. Table 6 includes two examples when analyzing spurious features in CelebA and one case of NICO++ . First, we trained a biased model with the original datasets. We already know that for the CelebA dataset, the spurious feature is gender (which is typically depicted by facial features instead of hair) and for the birds slice in NICO++, it is grass. We measured the overlap between the ground truth, which is the actual mask of the face (CelebA) / grass (NICO++), and the area highlighted by the VE methods, showing it clearly overlaps with the spurious features (in cyan).
Table 6: VEs are capable of identifying spurious correlations. We use IoU to measure the overlap between a segmentation representing the spurious feature and the mask we get from the VE.
Dataset Visual Features IoU
GradCAM GradCAM++ FullGrad ScoreCAM
CelebA (class blond) Hair $5.1\%$ $3.3\%$ $6.5\%$ $2.1\%$
Face $\mathbf{55.3\%}$ $\mathbf{61.5\%}$ $\mathbf{52.8\%}$ $\mathbf{57.6\%}$
NICO++95 (class bird) Rock $5.1\%$ $3.3\%$ $6.5\%$ $2.1\%$
Grass $\mathbf{70.4\%}$ $\mathbf{73.5\%}$ $\mathbf{69.8\%}$ $\mathbf{57.6\%}$
Dim lightning $2.3\%$ $2.1\%$ $1.8\%$ $2.2\%$
Outdoor $4.6\%$ $3.8\%$ $4.2\%$ $5.0\%$
Water $6.3\%$ $7.4\%$ $6.9\%$ $5.0\%$
Autumn $3.8\%$ $1.8\%$ $3.8\%$ $2.2\%$
0.A.4 Additional qualitative visualizations and results
0.A.4.1 Discovered Keywords by ViG-B2T
@@
0.A.4.2 Complete Tables of Keywords
Tables 7(b)-10(b) present a comparison between the keywords extracted by B2T and ViG-B2T. We can observe that ViG-B2T extracted keywords which are more relevant to the problem as they are related to spurious correlations, effectively highlighting biases in the dataset. Specifically, ViG-B2T extracted keywords which are related to gender (male or female), whereas B2T’s keywords only mention person without specifying gender, which is crucial information for bias discovery in this case.
@@ 0.A.4.2 Complete Tables of Keywords
0.A.4.3 Comparing biased and de-biased models using GradCAM
@@
0.A.4.4 Slices discovered by ViG-FACTS
@@
Note that the initial slices with low accuracy can be attributed to their association with bias-conflicting slices, representing minority groups in the dataset where the model performs poorly. Conversely, slices with higher accuracy typically correspond to bias-aligned groups, containing spurious visual features that the model relies on to make predictions. Finally, we observe that certain slices contain fewer images than others. This discrepancy arises because the FACTS method utilizes a mixture model to generate bias-aligned and bias-conflicting slices. Naturally, to evaluate this approach, the datasets we used contain ground truth for each slice. We extract the top-10 images based on the likelihood of each slice and assess the precision of these images. However, since these represent minority groups, there are instances where we have fewer than 10 images in a slice in the test data, leading to variations in the number of images.